OpenStudy (anonymous):

expand log(sqrt of x) and then log(a^2b)

OpenStudy (anonymous):

1.) log (sqrt (x)) = log x^(1/2) = 1/2 log(x) 2.) log (a^(2b)) = 2b log (a)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what is 2?

OpenStudy (anonymous):

a^ ?

OpenStudy (anonymous):

what 2?

OpenStudy (anonymous):

i cant read what was written for number 2 it doesnt make sense

OpenStudy (anonymous):

a^(2b) isnt that what you were asking for in the question?

OpenStudy (anonymous):

ya is that written as a to the 2b?

OpenStudy (anonymous):

yes it is. were you asking for log(a^2*b)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that would be log (a^2) + log (b) = 2log(a) = log(b)

OpenStudy (anonymous):

that should be 2log(a) + log(b)*

OpenStudy (anonymous):

lol so i need to rewrite the whole thing im confused. how does it look showin work

OpenStudy (anonymous):

is it log(a^(2b))??

OpenStudy (anonymous):

is it log(a^(2b))??

OpenStudy (anonymous):

is it log(a^(2b))??

OpenStudy (anonymous):

is it log(a^(2b))??

OpenStudy (anonymous):

it is log (a^2*b) = log (a^2) + log (b) from the multiplication rule = 2 log(a) + log(b) from the exponent rule

OpenStudy (anonymous):

Or you can just use the multiplication rule entirely: \[log(a^2b) \]\[= log(a\cdot a \cdot b) \]\[= log(a) + log(a) + log(b)\]\[=2log(a) + log(b)\]